# SLO-Aware Scheduling for Large Language Model Inferences

### v1

@@ 2.2. Related Work
In contrast, FastServe(Wu et al., 2024b) assigns priority to requests based on their prompt lengths and implements skip-join Multi-Level Feedback Queue scheduling along with iteration-level preemption to mitigate the head-of-line blocking inherent in the FCFS and run-to-completion approaches. However, FastServe is not specifically tailored for multi-SLO scenarios, as it bases request priorities solely on input lengths. Llumnix (Sun et al., 2024) is another work that takes the priority of requests into consideration. It highlighted the needs for new request rescheduling goals such as de-fragmentation and priorities. Nonetheless, it fails to provide an executable solution on how to determine the priority of individual requests based on their SLOs.
@@ 5. Evaluation
In this section, we evaluated the SLO-aware scheduler through a series of experiments. Initially, we showcase the overall performance enhancements over the current state-of-the-art vLLM framework. Extended experiments with varied framworks, models, and devices are also delivered and shown in Appendix due to the limited space. Next, we delve into the influence exerted by individual features. Finally, we test the scalability across instances.
Figure 7. Overall performance with Qwen2.5-7B in A100 evaluated in a mixed dataset of ShareGPT_Vicuna_unfiltered dataset and Python-Code-23k-ShareGPT dataset with varying request numbers and maximum batch sizes of (A) 1, (B) 2, and (C) 4. Note that results of exhaustive searching request number above 10 in (A), and 6 in (B) and (C) are not displayed, due to its execution inefficiency.
@@ 5.1. Experimental Settings
Testbed. The evaluation utilizes 2 NVIDIA V100 32GB GPUs by default. The NVIDIA V100 GPUs are connected over PCIe 3.0 x16, equipped with 516GB host memory and 2 Intel Xeon 6132 CPUs. Extended tests were also delivered delivered in 1 NVIDIA A800 GPU and shown in Appendix. The scalability tests were deployed to up to 8 V100 GPUs in the same server.
LLM models. The LLM model utilized in the experiment is Qwen2.5-7B (Yang et al., 2024), which has been widely used in industry and academia. One instance utilizes 2 NVIDIA V100 GPUs, and FP16 precision is used during the inference. Extended tests including Qwen2.5-7B in 1 A800, Qwen2.5-32B in 4 V100s, and Qwen2.5-32B in 1 A800 were also delivered.
@@
Baselines. The simulated-annealing SLO-aware scheduler operates as a dedicated scheduling component, necessitating its integration within a LLM service. For the experimental evaluation, we have adopted vLLM v0.6.0, a state-of-the-art LLM serving system renowned for its support of PageAttention and continuous batching, as our baseline platform. The simulated-annealing SLO-aware scheduler is implemented atop this vLLM system. Besides, we also used LMDeploy v0.7.2 to validate the effectiveness in varied framework, and the test results are shown in Appendix.
@@ 5.2. Overall Performance
request number 4 6 8 10
Simulated Annealing 0.00023 0.00032 0.00043 0.00048
Exhaustive Search 0.0012 0.039 88.2 287
Table 1. Algorithm overhead of simulated annealing-based priority mapping and exhaustive search.
@@
We compared the SLO attainment, the average latency, and their ratio (denoted as $G$) across varying request numbers and batch sizes. This evaluation was performed using the simulated-annealing SLO-aware scheduler (SA), an exhaustive-search counterpart, and vLLM which lacks SLO-awareness. Figure 7 presents the experimental outcomes for maximum batch sizes of 1 (illustrated in Figure 7 (A)), 2 (depicted in Figure 7 (B)), and 4 (shown in Figure 7 (C)). Results from exhaustive search for request numbers exceeding 8 may not be displayed, due to prohibitively long computing time. Owing to the variability in request sequences, the enhancement of the metric $G$ spans a range from 0.3% to 46.5%. Occasionally, the initial request execution order was sufficiently efficient with minimal waiting times, leading to only modest improvements from the simulated annealing process (as observed in the case with 10 requests and a batch size of 4 in Figure 7 (C)); or a slightly degradation (as seen with 6 requests and a batch size of 4 in Figure 7 (C)), as the result of variations in execution time. However, compared with vLLM, in most cases the simulated-annealing SLO-aware scheduler shows prominent improvement. It either elevated SLO attainment by up to 33.4% (observed in the case with 10 requests and the maximum batch size of 1 in Figure 7 (A)) or reduced average latency by up to 16.3% (seen in the case with 10 requests and the maximum batch size of 2 in Figure 7 (B)), achieving in an enhancement of the $G$ metric by up to 46.5% (when dealing with 10 requests and the maximum batch size of 1 in Figure 7 (A)).
Extended experiments in Appendix showed growths up to 5$\times$ in SLO attainment (with 40 requests and the maximum batch size of 1 in Qwen2.5-32B and LMDeploy at A800) and up to 31.6% (with 8 requests and the maximum batch size of 2 in Qwen2.5-7B and LMDeploy at A800) in average latency. A case with strict SLOs and a large request number (such as Qwen2.5-32B in 1 A800) produced a worse baseline and a larger search space, hence showed a more significant improvement in SLO attainment. On the contrary, the average latency improvement depended on more the baseline sequence randomness than the model or framework performance.
Furthermore, the simulate annealing-based priority mapping solution demonstrated performance with a maximum degradation of just 1.0% across the tested examples, compared to the exhaustive counterpart. We further compared the execution overhead of two priority algorithm with the maximum batch size of 1 and showed the results in Table 1. As request numbers increase, the exhaustive search solution’s execution time exhibits exponential growth, whereas that of the simulated annealing method demonstrates only a slight increase, which is attributed to the overhead of additional request management. This validates the efficiency of our simulated-annealing solution.
@@ 7. Acknowledgment
Appendix A Extended Experiments

### v2

@@ 2.2. Related Work
In contrast, FastServe(Wu et al., 2024b) assigns priority to requests based on their prompt lengths and implements skip-join Multi-Level Feedback Queue scheduling along with iteration-level preemption to mitigate the head-of-line blocking inherent in the FCFS and run-to-completion approaches. However, FastServe is not specifically tailored for multi-SLO scenarios, as it bases request priorities solely on input lengths. Llumnix (Sun et al., 2024) is another work that takes the priority of requests into consideration. It highlighted the needs for new request rescheduling goals such as de-fragmentation and priorities. Nonetheless, it fails to provide an executable solution on how to determine the priority of individual requests based on their SLOs. SOLA (Hong et al., [n.d.]) and Tempo (Zhang et al., 2025) aim to provide SLO-awareness for scheduling. However, multi-SLO scenarios are not considered.
@@ 5. Evaluation
In this section, we evaluated the SLO-aware scheduler through a series of experiments. Initially, we showcase the overall performance enhancements over varied framworks (vLLM and LMDeploy), models (Qwen2.5-7B and Qwen2.5-32B), and devices (NVIDIA V100 GPUs and NVIDIA A800 GPUs). Next, we delve into the influence exerted by individual features. Finally, we test the scalability across instances.
Figure 7. Overall performance with (A) Qwen2.5-7B in vLLM in V100, (B) Qwen2.5-7B in LMdeploy in V100, (C) Qwen2.5-32B in vLLM in V100, (D) Qwen2.5-32B in LMdeploy in V100, (E) Qwen2.5-7B in vLLM in A800, (F) Qwen2.5-7B in LMdeploy in A800, (G) Qwen2.5-32B in vLLM in A800, (H) Qwen2.5-32B in LMdeploy in A800, evaluated in a mixed dataset of ShareGPT_Vicuna_unfiltered dataset and Python-Code-23k-ShareGPT dataset with varying request numbers and maximum batch sizes.
@@ 5.1. Experimental Settings
Testbed. The evaluation utilizes 2 NVIDIA V100 32GB GPUs or 1 NVIDIA A800 80GB GPUs. The NVIDIA V100 GPUs are connected over PCIe 3.0 x16, equipped with 516GB host memory and 2 Intel Xeon 6132 CPUs. The NVIDIA A800 GPU is equipped with 1546GB and 2 Intel Xeon 6240 CPUs. The scalability tests were deployed to up to 8 V100 GPUs in the same server.
LLM models. The LLM models utilized in the experiment are Qwen2.5-7B and Qwen2.5-32B(Yang et al., 2024), which have been widely used in industry and academia. For Qwen2.5-7B, one instance utilizes 2 NVIDIA V100 GPUs, and FP16 precision is used during the inference. Experiments with Qwen2.5-7B in 1 A800, Qwen2.5-32B in 4 V100s, and Qwen2.5-32B in 1 A800 were also delivered.
@@
Baselines. The simulated-annealing SLO-aware scheduler operates as a dedicated scheduling component, necessitating its integration within a LLM service. For the experimental evaluation, we have adopted vLLM v0.6.0, a state-of-the-art LLM serving system renowned for its support of PageAttention and continuous batching, as our baseline platform. The simulated-annealing SLO-aware scheduler is implemented atop this vLLM system. Besides, we also used LMDeploy v0.7.2 to validate the effectiveness in varied framework.
@@ 5.2. Overall Performance
req num SA Exhaustive search
SLO attainment Latency (s) G (×$\times 10^{-5}$) Overhead (s) SLO attainment Latency (s) G (×$\times 10^{-5}$) Overhead (s)
4 1 10.29 9.72 0.00023 1 10.18 9.82 0.0012
6 0.83 11.44 7.26 0.00032 0.83 12.00 6.92 0.039
8 0.875 14.48 6.04 0.00043 0.875 14.50 6.03 88.2
10 0.4 27.61 1.45 0.00048 0.4 28.38 1.41 287
Table 1. Comparison of simulated annealing-based priority mapping and exhaustive search.
@@
We compared the SLO attainment, the average latency, and their ratio (denoted as $G$) across varying request numbers and batch sizes. This evaluation was performed using the simulated-annealing SLO-aware scheduler (SA), an exhaustive-search counterpart, and vLLM which lacks SLO-awareness. Figure 7 presents the experimental outcomes for (A) Qwen2.5-7B in vLLM in V100, (B) Qwen2.5-7B in LMdeploy in V100, (C) Qwen2.5-32B in vLLM in V100, (D) Qwen2.5-32B in LMdeploy in V100, (E) Qwen2.5-7B in vLLM in A800, (F) Qwen2.5-7B in LMdeploy in A800, (G) Qwen2.5-32B in vLLM in A800, (H) Qwen2.5-32B in LMdeploy in A800.
Owing to the variability in request sequences, the enhancement of the metric $G$ spans a range from 0 to 504%. Occasionally, the initial request execution order was sufficiently efficient with minimal waiting times, leading to only modest improvements from the simulated annealing process (as observed in the case with 40 requests and a batch size of 1 in Figure 7 (A)); or even no improvement (as seen with 8 requests and a batch size of 4 in Figure 7 (A)), as the result of variations in execution time. However, compared with vLLM and LMDeploy, in most cases the simulated-annealing SLO-aware scheduler shows prominent improvement. Among results with varied experiment settings showed growths up to 5$\times$ in SLO attainment (with 40 requests and the maximum batch size of 1 in Figure 7(D)) and up to 31.6% (with 8 requests and the maximum batch size of 2 in Figure 7(F)) in average latency. A case with strict SLOs and a large request number (as shown in Figure 7(G)-(H)) produced a worse baseline and a larger search space, hence showed a more significant improvement in SLO attainment. On the contrary, the average latency improvement depended on more the baseline sequence randomness than the model or framework performance. Besides, it can also be seen that sometimes the SLO attainment (seen in Figure 7 (B)) or the latency (seen in Figure 7 (A), (C), (F) and (H)) is sacrificed to achieve an overall higher $G$.
Furthermore, the simulate annealing-based priority mapping solution demonstrated performance with a maximum degradation of just 1.0% across the tested examples, compared to the exhaustive counterpart. We further compared the execution overhead of two priority algorithm with the maximum batch size of 1 and the request number across 4 to 10 and showed the results in Table LABEL:tab:overhead. Results for request numbers exceeding 10 were not recorded, due to prohibitively long computing time from exhaustive search. As request numbers increase, the exhaustive search solution’s execution time exhibits exponential growth, whereas that of the simulated annealing method demonstrates only a slight increase, which is attributed to the overhead of additional request management. This validates the efficiency of our simulated-annealing solution.
@@ 7. Acknowledgment
Appendix A Experiment Detailed Results
Figure 12. Qwen2.5-7B at 2 Nvidia V100 GPUs in vLLM